OpenStudy (anonymous):

In a certain county, the number of Charter Schools is ten less than twice the number of Alternative Schools. We know that there are 46 Charter Schools in the county. How many Alternative Schools are in the county? 28 56 18 40

OpenStudy (anonymous):

alternative school is x charter school is 2x-10 2x-10=x+46 2x-x=46+10 x=56

OpenStudy (anonymous):

therefore alternative schools is 56

OpenStudy (anonymous):